# Detecting Metadata-Related Bugs in Enterprise Applications

### v1

@@ Abstract.
When building enterprise applications (EAs) on Java frameworks (e.g., Spring), developers often configure application components via metadata (i.e., Java annotations and XML files). It is challenging for developers to correctly use metadata, because the usage rules can be complex and existing tools provide limited assistance. When developers misuse metadata, EAs become misconfigured, which defects can trigger erroneous runtime behaviors or introduce security vulnerabilities. To help developers correctly use metadata, this paper presents (1) RSL—a domain-specific language that domain experts can adopt to prescribe metadata checking rules, and (2) MeCheck—a tool that takes in RSL rules and EAs to check for rule violations.
With RSL, domain experts (e.g., developers of a Java framework) can specify metadata checking rules by defining content consistency among XML files, annotations, and Java code. Given such RSL rules and a program to scan, MeCheck interprets rules as cross-file static analyzers, which analyzers scan Java and/or XML files to gather information and look for consistency violations. For evaluation, we studied the Spring and JUnit documentation to manually define 15 rules, and created 2 datasets with 115 open-source EAs. The first dataset includes 45 EAs, and the ground truth of 45 manually injected bugs. The second dataset includes multiple versions of 70 EAs. We observed that MeCheck identified bugs in the first dataset with 100% precision, 96% recall, and 98% F-score. It reported 152 bugs in the second dataset, 49 of which bugs were already fixed by developers. Our evaluation shows that MeCheck helps ensure the correct usage of metadata.
@@ 1. Introduction
  • We designed a domain-specific language (DSL)—RSL—for domain experts to specify metadata-usage rules. Different from prior DSLs, RSL can express consistency relations among XML items, Java annotations, and Java code.
@@ 2. A Motivating Example
Figure 2. A bean object in XML has attribute init-method refer to a Java method in the corresponding class
As shown in Fig. 2, the exemplar XML file sets the attribute init-method of bean b to “myPostConstruct” (line 3). Ideally, EA developers should define a corresponding method in the exemplar Java class—myPostConstruct(). However, when EA developers fail to do so (see line 11), no existing compiler or static checker can reveal such errors of omission. As a result, developers can only observe the consequence of this metadata-related bug during program execution, and then manually diagnose root cause for the triggered runtime error org.springframework.beans.factory.BeanCreationException.
@@ 3.1. XML-Based Configuration
The multiple attributes of any element are separated by white spaces, and each attribute associates the attribute name with a string value. For simplicity, we use XML items to refer to XML elements and their attributes. For instance, Fig. 3 (a) shows an exemplar DD that defines two bean objects: infoMessage and messageRenderer (see lines 3–6). The second bean has a <constructor-arg> element (line 6), which references the first bean as the value of its attribute ref.
@@ 4.1. Statements
[…]
@@
[…]
@@ 4.3. Built-in Functions
XML-related function elementExists, getAttr, getAttrs, getElms, getXMLs, hasAttr
@@ 5. MeCheck
MeCheck conducts static program analysis, to check whether metadata is used correctly. As shown in Fig. 1, when being applied to a Java enterprise application, MeCheck (i) loads RSL rules, (ii) creates analyzers from those rules, and (iii) executes analyzers to detect bugs. This section focuses on steps (ii) and (iii) of the procedure.
@@ 5.2. Analyzer Executor
    Scoping: If an RSL rule define multiple variables using the same name, the executor should differenciate between the scopes of variables for correct static analysis.
@@ 5.2.1. Context-Aware Interpretation Algorithm
Input: $t$—root node of the rule/analyzer, and $P$—the given software application
@@ 0.4       $process(c,S,P)$
Input: $n$—statement node, $S$—stack for variable frames, and $P$—the given software application
@@ 0.3 $v\leftarrow$ variable name part from $n$’s header
[…]
0.5 foreach element $e$ in $container$ do
0.6       add {$T$, $v$, $e$} to $f$
@@ 0.8             $process(c,S,P)$
Input: $n$—statement node, $S$—stack for variable frames, and $P$—the given software application
@@ 0.3       $processFor(n,S,P)$
[…]
0.6       if $res$ is True then
@@ 0.18       $value\leftarrow evaluate($variable value part of $n,S,P)$
Input: $n$—statement node, $S$—stack for variable frames, and $P$—the given software application
@@
[…]
0.2 if $res$ = False then
0.3       $msg\leftarrow$ $evaluate($message part of the assert statement, $S,P)$
@@ 5.2.2. Stack and Frames
MeCheck adopts a stack $S$ to keep track of the execution context; it (1) pushes frames onto $S$, (2) pops frames from $S$, and (3) walks through frames to manage variables, their types, as well as their values. A frame is a dictionary, where the key is a variable name, and the value is a pair of variable type and variable name. MeCheck creates a frame before executing the entire analyzer, any ForStmt, IfStmt, or exists-clause; each created frame is then pushed onto stack to record any variable declared for or in the corresponding program structure. During execution, MeCheck refers to the stack to search for declared variables, query their data types, or obtain the values. After the execution of a ForStmt, IfStmt, exists-clause, or the entire analyzer, MeCheck pops a frame from the stack to discard all variables created by that program structure.
Input: $n$—expression node, $S$—stack for variable frames, and $P$—the given software application
@@
0.3       $res\leftarrow$ retrieve the value of $identifier$ from $S$ starting from top frame
0.4       $return$ $res$
@@
0.6       $return$ the literal’s value
@@
0.8       $return$ the function call’s return value
@@
[…]
@@
[…]
@@ 0.17       $container\leftarrow evaluate($container part from $n,S,P)$
0.18       foreach element $e$ in $container$ do
0.19             add {$T$, $v$, $e$} to $f$
@@ 0.21             $res\leftarrow$ $evaluate(logic,S,P)$
0.22             if $res$ = True then
0.23                   $return$ True
@@ 0.26       $return$ False
[…]
@@
[…]
@@
[…]
@@ 5.2.3. Performance Improvement
To optimize performance, we implemented a caching mechanism in MeCheck to eliminate redundant computation as much as possible. Specifically, we defined a global cache to map function calls to their return-values, for each software-under-analysis. For instance, as shown in Fig. 6 (a), getClasses() is mapped to the list of class items MeCheck creates based on Java parsing; getXMLs() is mapped to a list of XML items MeCheck creates based on XML parsing; getMethods("class-name") is mapped to a list of method items created for all methods declared by the Java class named “class-name”. We decided to cache a function call if that function extracts data from Java/XML files, or derives information by enumerating extracted data. With such a caching mechanism, for each project-under-analysis, MeCheck parses all Java (or XML) files only once if any of the analyzers needs to retrieve all code (or metadata) items; when a function is called with different parameter values, their results can be cached differently as those values are used as part of the key in cache. Furthermore, to avoid eagerly extracting all methods and fields from Java files when MeCheck is only interested in the class-level information, MeCheck conducts lazy initialization for both fields and methods members of ClassItem objects (see Fig. 6 (b)): either member is initialized only when any analyzer explicitly requests for the information of a particular class.
@@ 6.1. Effectiveness of RSL
To assess the expression power of RSL, we studied (1) the documentation of Spring (spr, 2021a)—the most popular application development framework for EAs in Java, and (2) documentation as well as tutorials on JUnit (jun, 2024b, a; how, 2024; sof, 2024)—the most popular framework for unit testing in Java. We manually distilled 15 metadata-usage constraints from those documents. As shown in Table 2, seven constraints are about the correspondence between XML and code ($r_{1}$$r_{7}$); six constraints involve annotations and code ($r8$$r13$); one constraint is on the correspondence between XML and annotation ($r14$); one constraint is relevant to XML, annotation, and code. These constraints are diverse in terms of the metadata/code items involved, the number of files scanned, and the consistency-checking logic. We successfully expressed constraint-checking rules in RSL, which fact evidences the great expression power of our domain-specific language. In the following experiments, we will leverage MeCheck to generate analyzers from these rules, and evaluate MeCheck’s effectiveness in identifying buggy EAs that violate any of the rules.
Finding 1 (Responding to RQ1): RSL has great expressiveness. It is capable of describing a variety of metadata-usage checking rules that involve XML, annotation, and/or code.
@@ 6.2.2. Metrics
$\vspace{-.5em}P=\frac{\mbox{\# of correctly reported bugs}}{\mbox{Total \# of bug reports}}$
@@
$\vspace{-.5em}R=\frac{\mbox{\# of correctly reported bugs}}{\mbox{Total \# of known bugs}}$
@@
$\vspace{-.5em}F=\frac{2\times P\times R}{P+R}$
All metrics mentioned above have values in [0%, 100%]; the higher the better. To compute P and R, we will intersect the set of bugs reported with the ground-truth bug set; any overlap between the two sets captures bugs correctly reported by MeCheck. Thus, the larger overlap there is, the better.
@@ 5 String beanClassFQN = getAttr(bean, class”);
Finding 2 (Responding to RQ2): MeCheck demonstrated great detection accuracy when being applied to our 45-project dataset with injected bugs. It found bugs with 100% precision, 96% recall, and 98% F-score.
@@ 6.3.2. Experiment and Results
Our results show that among the selected 70 software repositories, MeCheck reported in total 152 bugs in 21 projects. After manually inspecting the bug reports, we found 117 reports to be true positives, as they violate either $r_{1}$ (i.e., xml-path-check) or $r_{2}$ (i.e., bean-class-exists). In particular, 18 bugs were detected in the first commits of software repositories; 99 bugs were found in later commits, meaning that developers accidentally introduced the bugs when they revised software for maintenance. 115 of the 117 bugs were reported together with bugs of the same kind. The existence of multiple bugs in the same program version does not affect our evaluation results, as bugs are analyzed and reported independently.
Furthermore, we noticed that 49 of the 117 true positives were already fixed by developers. We found that out by checking later versions of the same projects and by observing revision of those buggy programs. The remaining 68 bugs had not been fixed yet, so we filed bug reports to contact developers and seek for their feedback. So far, we have not heard from developers for those bug reports. There are 35 false positives in the 152 bug reports, because the RegEx patterns we used to identify library classes do not cover all the libraries used by EAs.
@@ 14 ShcUtils (shc, 2024) 1 1 1 0
Finding 3 (Response to RQ3): MeCheck demonstrated great effectiveness and high performance when being applied to different versions of 70 real-world open-source EAs. It reported in total 152 bugs; 117 of the bugs are true positives, 49 of which have been already fixed by developers
@@ 7. Threats to Validity
Currently, MeCheck determines whether a given class $C$ is defined by a library dependency of EA using pattern matching. Namely, if the fully qualified name of $C$ matches a RegEx pattern predefined in MeCheck, it is a library class. However, our RegEx pattern set may not be comprehensive; some library classes may fail the matching process and get wrongly treated as non-library classes. One way to overcome this limitation is to eagerly scan the bytecode of all library dependencies, trying to find an exact match for $C$’s name. However, we decided not to perform such a heavyweight scanning for library classes because based on our experience, many open-source projects do not contain all JAR files for their library dependencies. Such a lack of dependency information will considerably limit the effectiveness of a heavyweight scanning and the applicability of MeCheck. In the future, we will investigate more advanced lightweight approaches to mitigate this issue.

### v2

@@ Abstract.
When building enterprise applications (EAs) on Java frameworks (e.g., Spring), developers often configure application components via metadata (i.e., Java annotations and XML files). It is challenging for developers to correctly use metadata, because the usage rules can be complex and existing tools provide limited assistance. When developers misuse metadata, EAs become misconfigured, which can trigger erroneous runtime behaviors or introduce security vulnerabilities. To help developers correctly use metadata, this paper presents (1) RSL—a domain-specific language that domain experts can adopt to prescribe metadata checking rules, and (2) MeCheck—a tool that takes in RSL rules and EAs to check for rule violations.
With RSL, domain experts (e.g., owner developers of a Java framework) can specify metadata checking rules by defining content consistency among XML files, annotations, and Java code. Given such RSL rules and a program to scan, MeCheck interprets rules as cross-file static analyzers that scan Java and/or XML files to gather information and look for consistency violations. For evaluation, we studied the Spring and JUnit documentation to manually define 15 rules, and created 2 datasets with 115 open-source EAs. The first dataset includes 45 EAs, and the ground truth of 45 manually injected bugs. The second dataset includes multiple versions of 70 EAs. We observed that MeCheck identified bugs in the first dataset with 100% precision, 96% recall, and 98% F-score. It reported 152 bugs in the second dataset, 49 of which were already fixed by developers. Our evaluation shows that MeCheck helps ensure the correct usage of metadata.
@@ 1. Introduction
  • We designed a domain-specific language (DSL)—RSL—for domain experts to specify metadata-usage rules. Unlike prior DSLs, RSL can express consistency relations among XML items, Java annotations, and Java code.
@@ 2. A Motivating Example
Figure 2. A bean in XML has init-method refer to a nonexistent method in the corresponding class
As shown in Fig. 2, the exemplar XML file sets the attribute init-method of bean b to “myPostConstruct” (line 3). Ideally, EA developers should define a corresponding method in the exemplar Java class—myPostConstruct(). However, when EA developers fail to do so (see line 11), no existing compiler or static checker can reveal such omissions. As a result, developers can only observe the consequence of this metadata-related bug during program execution, and then manually diagnose root cause for the triggered runtime error org.springframework.beans.factory.BeanCreationException.
@@ 3.1. XML-Based Configuration
The multiple attributes of any element are separated by white spaces, and each attribute associates the attribute name with a string value. For simplicity, we use XML items to refer to XML elements and their attributes. For instance, Fig. 3 (a) shows an exemplar DD that defines two bean objects: infoMessage and messageRenderer (see lines 3–6). The second bean has a <constructor-arg> element (line 5), which references the first bean as the value of its attribute ref.
@@ 4.1. Statements
$\textrm{Specification}:=$ $\;\textbf{Rule}\ {\tt Id}\ \textrm{Body}$
$\textrm{Body}:=$ $^{\prime}\{^{\prime}\textrm{Stmt}\ \textrm{Stmt}*^{\prime}\}^{\prime}$
$\textrm{Stmt}:=$ $\;\textrm{ForStmt}\ |\ \textrm{IfStmt}\ |\ \textrm{AssertStmt}\ |\ \textrm{DeclStmt}\ ^{\prime};^{\prime}$
$\textrm{ForStmt}:=$ $\;\textbf{for}\ ^{\prime}(^{\prime}\ \textrm{Type}\ {\tt Id}\ \textrm{in}\ \textrm{Exp}\ ^{\prime})^{\prime}\ \textrm{Body}$
$\textrm{IfStmt}:=$ $\;\textbf{if}\ ^{\prime}(^{\prime}\ \textrm{Exp}\ ^{\prime})^{\prime}\ \textrm{Body}$
$\textrm{AssertStmt}:=$ $\;\textbf{assert}\ ^{\prime}(^{\prime}\ \textrm{Exp}\ ^{\prime})^{\prime}\ {}^{\prime}\{^{\prime}\ \textrm{MsgStmt}\ ^{\prime};^{\prime}\ {}^{\prime}\}^{\prime}$
$\textrm{MsgStmt}:=$ $\;\textbf{msg}\ ^{\prime}(^{\prime}\ {}^{\prime},^{\prime}\ \textrm{SimExp}\ (^{\prime},^{\prime}\ \textrm{SimExp})*\ ^{\prime})^{\prime}$
$\textrm{DeclStmt}:=$ $\;\textrm{Type}\ {\tt Id}\ ^{\prime}=^{\prime}\ \textrm{Exp}$
$\textrm{Exp}:=$ $\;\textrm{SimExp}\ |\ \textrm{SimExp}\ \textbf{AND}\ \textrm{Exp}\ |\ \textrm{SimExp}\ \textbf{OR}\ \textrm{Exp}\ |\textbf{NOT}\ \textrm{Exp}$
$\textrm{SimExp}:=$ $\;{\tt Id}\ |\ {\tt Lit}\ |\ \textrm{FunctionCall}\ |\ ^{\prime}(^{\prime}\ \textrm{Exp}\ ^{\prime})^{\prime}\ |\ \textrm{FunctionCall}\ ^{\prime}==^{\prime}\textrm{SimExp}\ |\ \textbf{exists}\ ^{\prime}(^{\prime}\ {\tt Type}\ {\tt Id}\ \textbf{in}\ \textrm{Exp}\ ^{\prime})^{\prime}\ {}^{\prime}(^{\prime}\ \textrm{Exp}\ ^{\prime})^{\prime}$
@@
$\textrm{FunctionCall}:=$ $\;{\tt Id}\ ^{\prime}(^{\prime}\ \textrm{Params}\ ^{\prime})^{\prime}$
$\textrm{Params}:=$ $\;\textrm{SimExp}\ (^{\prime},^{\prime}\ \textrm{SimExp})*$
@@ 4.3. Built-in Functions
XML-related functions elementExists, getAttr, getAttrs, getElms, getXMLs, hasAttr
@@ 5. MeCheck
MeCheck conducts static program analysis, to check whether metadata is used correctly. As shown in Fig. 1, when applied to a Java enterprise application, MeCheck (i) loads RSL rules, (ii) creates analyzers from those rules, and (iii) executes them to detect bugs. This section focuses on steps (ii) and (iii) of this process.
@@ 5.2. Analyzer Executor
    Scoping: If an RSL rule defines multiple variables using the same name, the executor should differentiate between the scopes of variables for correct static analysis.
@@ 5.2.1. Context-Aware Interpretation Algorithm
Input: tt—root node of the rule/analyzer, and PP—the given software application
@@ 0.4  $process(c,S,P)$
Input: nn—statement node, SS—stack for variable frames, and PP—the given software application
@@ 0.3 $v\leftarrow$ variable name part from $n$’s header
0.4 $container\leftarrow evaluate($container part from $n$’s header,S,P),S,P)
0.5 foreach element $e$ in containercontainer do
0.6  add {$T$, $v$, ee} to $f$
@@ 0.8     $process(c,S,P)$
Input: nn—statement node, SS—stack for variable frames, and PP—the given software application
@@ 0.3  $processFor(n,S,P)$
0.5  $res\leftarrow evaluate(n$’s expression,S,P),S,P)
0.6  if resres is True then
@@ 0.18  $value\leftarrow evaluate($variable value part of $n,S,P)$
Input: nn—statement node, SS—stack for variable frames, and PP—the given software application
@@
0.1 $res\leftarrow$ $evaluate(n$’s expression,S,P),S,P)
0.2 if resres = False then
0.3  $msg\leftarrow$ evaluate(evaluate(message part of the assert statement, $S,P)$
@@ 5.2.2. Stack and Frames
MeCheck adopts a stack $S$ to keep track of the execution context; it (1) pushes frames onto $S$, (2) pops frames from $S$, and (3) walks through frames to manage variables, their types, as well as their values. A frame is a dictionary, where the key is a variable name, and the value is a pair of data type and variable value. MeCheck creates a frame before executing the entire analyzer, any ForStmt, IfStmt, or exists-clause; each created frame is then pushed onto stack to record any variable declared for or in the corresponding program structure. During execution, MeCheck refers to the stack to search for declared variables, query their data types, or obtain the values. After the execution of a ForStmt, IfStmt, exists-clause, or the entire analyzer, MeCheck pops a frame from the stack to discard all variables created by that program structure.
Input: nn—expression node, SS—stack for variable frames, and PP—the given software application
@@
0.3  $res\leftarrow$ retrieve the value of identifieridentifier from $S$ starting from top frame
0.4  returnreturn resres
@@ 0.3  $res\leftarrow$ retrieve the value of identifieridentifier from $S$ starting from top frame
0.6  returnreturn the literal’s value
@@
0.8  returnreturn the function call’s return value
@@
0.10  returnreturn evaluate(evaluate(the expression inside the parenthesis,S,P),S,P)
@@
0.12  returnreturn evaluate(evaluate(left-side of the expression,S,P),S,P) == evaluate(evaluate(right-side of the expression,S,P),S,P)
@@ 0.17  $container\leftarrow evaluate($container part from $n,S,P)$
0.18  foreach element $e$ in containercontainer do
0.19     add {$T$, $v$, ee} to $f$
@@ 0.21     $res\leftarrow$ $evaluate(logic,S,P)$
0.22     if resres = True then
0.23        returnreturn True
@@
0.28  returnreturn evaluate(evaluate(left-side of the expression,S,P),S,P) && evaluate(evaluate(right-side of the expression,S,P),S,P)
@@
0.30  returnreturn evaluate(evaluate(left-side of the expression,S,P),S,P) —— evaluate(evaluate(right-side of the expression,S,P),S,P)
@@
0.32  returnreturn  evaluate(evaluate(the expression after NOT operator,S,P),S,P)
@@ 5.2.3. Performance Improvement
To optimize performance, we implemented a caching mechanism in MeCheck to eliminate redundant computation as much as possible. Specifically, we defined a global cache to map function calls to their return-values, for each software-under-analysis. For instance, as shown in Fig. 6 (a), getClasses() is mapped to the list of class items MeCheck creates based on Java parsing; getXMLs() is mapped to a list of XML items MeCheck creates based on XML parsing; getMethods("class-name") is mapped to a list of method items created for all methods declared by the Java class named “class-name”. We decided to cache a function call if that function extracts data from Java/XML files, or derives information by enumerating extracted data. With such a caching mechanism, for each project-under-analysis, MeCheck parses all Java (or XML) files only once if any of the analyzers needs to retrieve all code (or metadata) items; when a function is called with different parameter values, their results can be cached differently as those values are used as part of the cache key. Furthermore, to avoid eagerly extracting all methods and fields from Java files when MeCheck is only interested in the class-level information, MeCheck conducts lazy initialization for both fields and methods members of ClassItem objects (see Fig. 6 (b)): either member is initialized only when any analyzer explicitly requests for the information of a particular class.
@@ 6.1. Effectiveness of RSL
To assess the expression power of RSL, we studied (1) the documentation of Spring (spr, 2021a)—the most popular application development framework for EAs in Java, and (2) documentation as well as tutorials on JUnit (jun, 2024b, a; how, 2024; sof, 2024)—the most popular framework for unit testing in Java. We manually distilled 15 metadata-usage constraints from those documents. As shown in Table 2, seven constraints are about the correspondence between XML and code ($r_{1}$$r_{7}$); six constraints involve annotations and code (r8r8r13r13); one constraint is on the correspondence between XML and annotation ($r14$); one constraint is relevant to XML, annotation, and code. These constraints are diverse in terms of the metadata/code items involved, the number of files scanned, and the consistency-checking logic. We successfully expressed constraint-checking rules in RSL, which fact evidences the great expression power of our domain-specific language. In the following experiments, we will leverage MeCheck to generate analyzers from these rules, and evaluate MeCheck’s effectiveness in identifying buggy EAs that violate any of the rules.
Finding 1 (Answer to RQ1): RSL has great expressiveness. It is capable of describing a variety of metadata-usage checking rules that involve XML, annotation, and/or code.
@@ 6.2.2. Metrics
[…]
@@
[…]
@@
$F=\frac{2\times P\times R}{P+R}$
All metrics mentioned above have values in [0%, 100%]; the higher, the better. To compute P and R, we will intersect the set of bugs reported with the ground-truth bug set; any overlap between the two sets captures bugs correctly reported by MeCheck. Thus, the larger overlap there is, the better.
@@ 5 String beanClassFQN = getAttr(bean, class”);
Finding 2 (Answer to RQ2): MeCheck demonstrated great detection accuracy when being applied to our 45-project dataset with injected bugs. It found bugs with 100% precision, 96% recall, and 98% F-score.
@@ 6.3.2. Experiment and Results
Our results show that among the selected 70 software repositories, MeCheck reported in total 152 bugs in 21 projects. After manually inspecting the bug reports, we found 117 reports to be true positives, as they violate either $r_{1}$ (i.e., xml-path-check) or $r_{2}$ (i.e., bean-class-exists). In particular, 18 bugs were detected in the first commits of software repositories; 99 bugs were found in later commits, meaning that developers accidentally introduced the bugs when they revised software for maintenance. 115 of the 117 bugs were reported together with bugs of the same kind. The existence of multiple bugs in the same program version does not affect our evaluation results, as bugs are analyzed and reported independently. Furthermore, we noticed that 49 of the 117 true positives were already fixed by developers. We found that out by checking later versions of the same projects and by observing revision of those buggy programs. The remaining 68 bugs had not been fixed yet, so we filed bug reports to contact developers and seek for their feedback. So far, we have not heard from developers for those bug reports. There are 35 false positives in the 152 bug reports, because the RegEx patterns we used to identify library classes do not cover all the libraries used by EAs.
@@ 14 ShcUtils (shc, 2024) 1 1 1 0
Finding 3 (Answer to RQ3): MeCheck demonstrated great effectiveness and high performance when being applied to different versions of 70 real-world open-source EAs. It reported in total 152 bugs; 117 of the bugs are true positives, 49 of which have been already fixed by developers
@@ 7. Threats to Validity
Currently, MeCheck determines whether a given class $C$ is defined by a library dependency of EA using pattern matching. Namely, if the fully qualified name of $C$ matches a RegEx pattern predefined in MeCheck, it is considered a library class. However, our RegEx pattern set may not be comprehensive; some library classes may fail the matching process and get wrongly treated as non-library classes. One way to overcome this limitation is to eagerly scan the bytecode of all library dependencies, trying to find an exact match for $C$’s name. However, we decided not to perform such a heavyweight scanning for library classes because based on our experience, many open-source projects do not contain all JAR files for their library dependencies. Such a lack of dependency information will considerably limit the effectiveness of a heavyweight scanning and the applicability of MeCheck. In the future, we will investigate more advanced lightweight approaches to mitigate this issue.